# Exceptional-point-controlled mode interaction in three-dimensional microcavities represented by generalized Husimi functions

### v1

@@ Abstract
Non-Hermitian photonics has attracted significant interest and influences several key areas such as optical metamaterials, laser physics, and nonlinear optics. While non-Hermitian effects have been widely addressed in two-dimensional systems, we focus on realistic three-dimensional devices. To this end we generalize established phase space methods from mesoscopic optics and introduce Husimi functions for three-dimensional systems that deepen the insight and access to the mode morphology and their dynamics. We illustrate that four-dimensional Husimi functions can be represented using a specific projection in two dimensions and illustrate it for (conical) cylindrical cavities. The non-Hermitian character of the intrinsically open photonic systems is in particular revealed when examining the TE and TM polarization character of the resonance modes. Unlike the 2D case, polarization is not conserved in three-dimensional cavities, and we use generalized Husimi function to represent the interaction of polarization modes. We find their dynamics to be ruled by a network of exceptional points in the parameter space spanned by the refractive index and the cavity geometry tilt angle. This approach not only enhances our understanding of cavity modes but also aids in the design of more efficient photonic devices and systems.
@@ I Introduction
Optical microcavities [1, 2, 3] have proven to be fascinating and application relevant model system for several decades. Their intrinsic openness – light can enter and leave the cavity by refraction or evanescent coupling – is the source of fascinating properties and devices ranging from microlasers to sensors. These open optical microcavities of various geometries have considerably enriched the field of quantum chaos beyond the hard-wall billiard systems typically studied in nonlinear dynamics [4] . Their description in both real and phase space, has contributed significantly to their understanding in terms of ray-wave correspondence [5, 6, 7], leading to the development and application of deformed microcavity lasers with directional emission [8, 9, 10, 11, 12]. Other notable applications of photonic microsystems include optical gyroscopes [13, 14], cavity QED studies [15, 16, 17] and single particle detection [18, 19, 20].
@@
Moreover, the openness of such mesoscopic optical cavities yields the remarkable feature of non-Hermitian physics since all resonances are characterized by a complex frequency $\omega$ (with the quality $Q$ factor being related to its imaginary part by $Q=-\mathrm{Re}\,(\omega)/2\mathrm{Im}\,(\omega)$). The resulting non-Hermitian effects have been studied in depth in the past years in a variety of systems such as microlasers [21, 22, 23, 24, 25, 26], sensors [27, 28, 29, 30, 31, 32, 33], encryption [34], topological devices [35, 36, 37, 38, 39] and optical gyroscopes [13, 14]. Non-Hermitian properties are not limited to optical systems. They can be realized in a variety of fields, such as acoustic [40, 41], optomechanical [42], magnetic [43] and plasmonic devices [44], as well as integrated electrical circuits [45].
@@ II Cylindrical waveguides and cross-sectional Husimi functions
[…]
@@
$[\partial_{x}^{2}+\partial_{y}^{2}+\tilde{n}^{2}(\bm{r})k^{2}]\,\psi(x,y)=0$ (1)
@@
with angular momentum quantum number $m\in\mathbb{Z}$ and the Bessel (Hankel) function $J_{m}(x)$ ($H_{m}^{(1)}(x)$) of the first kind. Note that resonances with ±mplus-or-$\pm m$ are doubly degenerate. The values for $c_{1,2}$ are obtained by considering the appropriate boundary conditions. This introduces a quantization in $r$, which is described by the quantum number $\rho\in\mathbb{N}_{0}$.
Two example solutions are shown in Fig. 1a and b. Fig. 1a shows the electric field intensity $|\psi|^{2}$from Eq. (3) for a TM mode with $(\rho,m)=(0,5)$. Fig. 1b shows $|\psi|^{2}$ for the corresponding TE mode. These morphologies are also called whispering-gallery modes, as they travel along the cavity surface where they are confined by total internal reflection. Note that the wave function in Fig. 1 is calculated using a 3D cavity, however the solutions for the out-of-plane component are equivalent to the 2D case [63, 64].
Phase space methods are a well-developed tool to characterize optical systems with ray and wave dynamics. The state of a particle in a two-dimensional billiard system can be described by four generalized coordinates $(\bm{q},\bm{p})$ with $\bm{q},\bm{p}\in\mathbb{R}^{2}$ in a four-dimensional phase space. In Hamiltonian systems, conservation of energy confines a particle to a 3D manifold. In addition, a particle trajectory is completely described by considering its reflection points at the interface boundary, with location $s$ and angle of incidence $\chi$. Consequently, a 2D phase space is sufficient to describe the dynamics of a 2D system. An example for such a projection is the well-known Poincaré surface of section for billiards/ray optics [50, 6, 51, 65]. The Husimi function is the counterpart of the trajectory trace for wave dynamical systems and can be computed based on the wave function and its derivative [52, 53, 56, 66, 7], see Eq. (10) below.
The 2D Husimi function is based on $\psi(s)$ and $\psi^{\prime}(s)$ where $s$ is the coordinate along the system boundary and indicates the normal derivative. The other phase space coordinate is $\sin\chi$, the sine of the (classical) angle of incidence that is related to angular momentum $m$ by $\sin\chi=m/n\mathrm{Re}\,(\Omega)$ with $\mathrm{Re}\,(\Omega)=kR_{0}$. The Husimi function $H(s,\sin\chi)$ for the wave function in Fig. 1a(b) is shown in Fig. 1c(d). The mode is a radially symmetric whispering-gallery mode, and the angular momentum is thus conserved. Note that this is the simplest case, as for deformed cavities more intricate wave function (and thus phase space) emerges.
@@ III.1 Cylindrical cavities of finite height
$[\Delta+\tilde{n}^{2}(\bm{r})k^{2}]\begin{Bmatrix}\bm{E}(\bm{r})\\ \bm{B}(\bm{r})\end{Bmatrix}=\bm{0}$ (4)
@@
[…]
@@
Applying the appropriate boundary conditions leads to a quantization in the $z$-direction, described by the quantum number $\eta\in\mathbb{N}_{0}$. So in general, modes are identified by their polarization and the three quantum numbers $(\rho,m,\eta)$. Examples for confinement in the $z$ direction for TM (TE) modes are shown in Fig. 1e (f). The TM mode is more confined in $z$ direction due to the discontinuity in $E_{z}$ described in Eq. (5).
@@ III.2 Extension of the Husimi formalism to three-dimensional cavities
For realistic 3D systems that we want to describe now, the new generalized coordinates $(\bm{q},\bm{p})$ with $\bm{q},\bm{p}\in\mathbb{R}^{3}$ result in a 6D phase space. It reduces to four dimensions after considering conservation of energy and projection to a surface of section that we choose to be the cylinder mantle here.
We map the wave function onto the cylinder mantle using $\psi(\varphi,z)\rightarrow\psi(s,z)$, where $s=\varphi R_{0}$ is the arclength along the cavity surface. The resulting boundary conditions are periodic in $s$ with $\psi(s,z)=\psi(s\pm 2\pi R_{0},z)$. Here, we focus on the mantle as the main surface contribution, although top and bottom surfaces could be included straightforwardly when needed, e.g., for flat cavities.
For the phase space coordinates we follow the conventional notation of the Husimi function formalism [52]. For 3D cavities the Husimi function $H=H(\bm{q},\bm{p})$ depends on the spatial coordinates $\bm{q}=(s,z)$ and on the momentum coordinates $\bm{p}=(nk\sin\chi_{xy},nk\sin\chi_{z})$, see Fig. 2a. The momentum coordinates can also be written in polar representation $(nk\sin\chi_{0},\phi_{xy})$ with total momentum in the surface (mantle) plane $n^{2}k^{2}\sin^{2}\chi_{0}=|\bm{p}|^{2}=n^{2}k^{2}(\sin^{2}\chi_{xy}+\sin^{2}\chi_{z})$ and angle $\phi_{xy}$ to the $xy$-plane where $\tan\phi_{xy}=\sin\chi_{z}/\sin\chi_{xy}$.
@@
$\xi(\bm{q}^{\prime};\bm{q},\bm{p})=\frac{1}{\sqrt{\pi\sigma^{2}}}\times\\ \exp\left[-|\bm{q}^{\prime}-\bm{q}|^{2}/2\sigma^{2}-i\,\bm{p}\cdot(\bm{q}^{\prime}-\bm{q})\right]$ (8)
with variance $\sigma^{2}$. We then define its overlap with the wave function $\psi$ at the boundary as
@@
with angular momentum weight $F(p)=\sqrt{n}(1-\sin^{2}\chi_{0})^{1/4}$.
To illustrate this generalized Husimi function in the new phase space spanned by $\sin\chi_{xy}$ and $\sin\chi_{z}$, it is helpful to consider a simple test system: A wave function $\psi$ is given by a linear combination of plane waves, which propagate along a two-dimensional surface as shown in Fig. 2b-c. The Husimi function is calculated for a representative, fixed position $\bm{q}$ while the momentum $\bm{p}$ is varied. A single plane wave can be identified trivially in real space and gives a single peak in phase space in Fig. 2d. For two (or more) overlapping plane waves in Fig. 2c the field function can become arbitrarily complex, making it impossible to distinguish components in real space. The phase space recovers the constituent plane waves and their parameters as shown in Fig. 2e.
@@ III.3 Husimi functions for cylindrical cavities
[…]
@@
The Husimi function for the $(\rho,m,\eta)=(0,5,0)$-TM(TE) modes with $h/R=0.75$ are shown in Fig. 1e (f). The phase space figures are calculated by using a constant coherent state position $\bm{q}=\mathrm{const}$ and varying the momentum coordinates $\bm{p}$. This gives a circular border for the parameter space from the constraint $\sin^{2}\chi_{xy}+\sin^{2}\chi_{z}<1$. Due to the rotational symmetry, only coherent state positions with angles $\varphi\in[0,\pi/2m]$ need to be considered. Given the mirror symmetry at $z=0$, we focus on coherent states with $z\in[0,-h/2]$ without loss of generality.
@@
A sharp peak is present at $\sin\chi_{xy}=\pm m/n\,\mathrm{Re}\,\Omega$ for all values of $\sin\chi_{z}$. This is analogous to the result of the 2D cavity in Figs. 1c and f. The phase space is mirrored along $\sin\chi_{xy}=0$ because the standing wave considered here is a superposition of two traveling waves with ±mplus-or-$\pm m$. In addition, the existence of the geometry symmetry plane, $z=0$, results in a phase space mirror symmetry with axis $\sin\chi_{z}=0$. For the TE mode, the intensity maximum is closer to the origin which reflects its shorter lifetime since a larger part of the phase space intensity is located in the leaky region.
@@
Another degree of freedom we need to take into account is the localization of the coherent state, given by its standard deviation $\sigma$ in Eq. (8). Fig. 3 compares the phase spaces for two values of $\sigma$ for various points $\bm{q}_{i}$ on the cylinder surface. Figures 3b,d,f and h show the results for strong localization with $\sigma/R_{0}=0.32$, corresponding to a full-width at half-maximum of $\lambda/n$, where $\lambda$ is the vacuum wavelength. The intensity distribution is different between $\bm{q}_{1}$ and $\bm{q}_{2}$ in Figs. 3b and d. $\bm{q}_{1}$ is located on an intensity maximum, leading to $H(\bm{q}=\bm{q}_{1},\bm{p}=\bm{0})>0$. The nodal line at $\bm{q}_{2}$ results in $H(\bm{q}=\bm{q}_{2},\bm{p}=\bm{0})=0$, since the fields of two neighboring intensity maxima average to zero due to their different signs. The Husimi functions in Figs. 3f and h are taken at points closer to the cavity boundary. The intensity at $\bm{q}=\bm{q}_{3},\bm{q}_{4}$ is significantly lower than at $\bm{q}=\bm{q}_{1},\bm{q}_{2}$, which is due to the lower overall field intensity at the cylinder end faces. Spatial dependence is very pronounced for localized coherent states and is lost for the weakly localized state with $\sigma=0.95$, see Figs. 3c,e,g and i.
@@
Fig. 4 explores the dependence of the Husimi function on the quantization in $z$ for a cylinder with $h/R=5$. It shows the behavior of the Husimi function with $\sigma/R_{0}=0.95$ for $\eta=0,3,5$ for TM modes with $(\rho,m)=(0,6)$. The wave functions are now described by sharply localized intensity maxima, in contrast to the previously spread out maxima in Fig. 3. When $\eta$ is increased the angle $\varphi_{xy}$ of the maximum increases. This is related to the distribution of wavenumber components in the in- and out-of-plane components. The mode performs whispering-gallery type motion in $xy$ and a ”bouncing ball” type motion in $z$. The rays can be thought of as angled at $\tan\phi_{xy}=k_{z}/k_{xy}$ (compare Fig. 2a). Increasing $\eta$ results in more energy being stored in the bouncing ball motion, thereby increasing $\sin\chi_{z}$ of the intensity maximum. The symmetry in the $xy$-plane results in two symmetry-related signatures at ±$\pm\sin\chi_{z}$. The wave function on the cylinder mantle displays a ”checkerboard” pattern, which can be recovered by the superposition of two plane waves at angles ±$\pm\varphi_{xy}$ (compare Figs. 2c,e). For TE modes the behavior is similar, which is demonstrated in Appendix A.

### v2

@@ Abstract
Non-Hermitian photonics has attracted significant interest and influences several key areas such as optical metamaterials, laser physics, and nonlinear optics. While non-Hermitian effects have been widely addressed in two-dimensional systems, we focus on realistic three-dimensional devices. To this end we generalize established phase space methods from mesoscopic optics and introduce Husimi functions for three-dimensional systems that deepen the insight and access to the mode morphology and their dynamics. We illustrate that four-dimensional Husimi functions can be represented using a specific projection on two dimensions and illustrate it for (conical) cylindrical cavities. The non-Hermitian character of the intrinsically open photonic systems is in particular revealed when examining the TE and TM polarization character of the resonance modes. Unlike the 2D case, polarization is not conserved in three-dimensional cavities, and we use generalized Husimi function to represent the interaction of polarization modes. We find their dynamics to be ruled by a network of exceptional points in the parameter space spanned by the refractive index and the cavity geometry tilt angle. This approach not only enhances our understanding of cavity modes but also aids in the design of more efficient photonic devices and systems.
Published in: Physical Review A 112, 033528 (2025)
DOI: 10.1103/2wrn-z5fg
@@ I Introduction
Optical microcavities [1, 2, 3] have proven to be a fascinating and application relevant model system for several decades. Their intrinsic openness – light can enter and leave the cavity by refraction or evanescent coupling – is the source of fascinating properties and devices ranging from microlasers to sensors. These open optical microcavities of various geometries have considerably enriched the field of quantum chaos beyond the hard-wall billiard systems typically studied in nonlinear dynamics [4]. Their description in both real and phase space has contributed significantly to their understanding in terms of ray-wave correspondence [5, 6, 7], leading to the development and application of deformed microcavity lasers with directional emission [8, 9, 10, 11, 12]. Other notable applications of photonic microsystems include cavity QED studies [13, 14, 15] and single particle detection [16, 17, 18].
@@
Moreover, the openness of such mesoscopic optical cavities yields the remarkable feature of non-Hermitian physics since all resonances are characterized by a complex frequency $\omega$ (with the quality $Q$ factor being related to its imaginary part by $Q=-\mathrm{Re}\,(\omega)/2\mathrm{Im}\,(\omega)$) [19, 20, 21] . The resulting non-Hermitian effects have been studied in depth in the past years in a variety of systems such as microlasers [22, 23, 24, 25, 26, 27], sensors [28, 29, 30, 31, 32, 33, 34], encryption [35], topological devices [36, 37, 38, 39, 40] and optical gyroscopes [41, 42]. In addition to optical systems, non-Hermitian properties can also arise in acoustic [43, 44], optomechanical [45], magnetic [46] and plasmonic devices [47], as well as integrated electrical circuits [48].
@@ II Cylindrical waveguides and cross-sectional Husimi functions
Figure 1: (a,b,e,f) Electric (red) and magnetic (blue) field intensities of cylinder modes with quantum numbers $(\rho,m,\eta)=(0,5,0)$ and refractive index $n=2.5$. The mode in (a,e) is TM-polarized with $\Omega_{\mathrm{TM}}=3.361-i\,0.010$. The mode in (b,f) is TE polarized mode with $\Omega_{\mathrm{TE}}=3.578-i\,0.014$. (e,f) Incident Husimi functions for the cross-sectional $(x,y)$ plane with (c) $\Omega_{\mathrm{TM}}^{\mathrm{2D}}=2.951-i\,0.010$ and (d) $\Omega_{\mathrm{TE}}^{\mathrm{2D}}=3.342-i\,0.017$. The dashed lines mark minimum angles for total internal reflection.
@@
$\partial_{x}^{2}+\partial_{y}^{2}+\tilde{n}^{2}(\bm{r})k^{2}]\,\psi(x,y)=0$ (1)
@@
with angular momentum quantum number $m\in\mathbb{Z}$ and the Bessel (Hankel) function $J_{m}(x)$ ($H_{m}^{(1)}(x)$) of the first kind. Note that resonances with $\pm m$ are doubly degenerate. The values for $c_{1,2}$ are obtained by considering the appropriate boundary conditions. This introduces a quantization in $r$, which is described by the quantum number $\rho\in\mathbb{N}_{0}$.
Two example solutions are shown in Fig. 1(a) and (b). Fig. 1(a) shows the electric field intensity $|\psi|^{2}$ from Eq. (3) for a TM mode with $(\rho,m)=(0,5)$. Fig. 1(b) shows $|\psi|^{2}$ for the corresponding TE mode. These morphologies are also called whispering-gallery modes, as they travel along the cavity surface where they are confined by total internal reflection. Note that the wave function in Fig. 1 is calculated using a 3D cavity, however the solutions for the out-of-plane component are equivalent to the 2D case [66, 67].
Phase space methods are a well-developed tool to characterize optical systems with ray and wave dynamics. The state of a particle in a two-dimensional billiard system can be described by four generalized coordinates $(\bm{q},\bm{p})$ with $\bm{q},\bm{p}\in\mathbb{R}^{2}$ in a four-dimensional phase space. In Hamiltonian systems, conservation of energy confines a particle to a 3D manifold. In addition, a particle trajectory is completely described by considering its reflection points at the interface boundary, with location $s$ and angle of incidence $\chi$. Consequently, a 2D phase space is sufficient to describe the dynamics of a 2D system. An example for such a projection is the well-known Poincaré surface of section for billiards and ray optics [53, 6, 54, 68]. The Husimi function is the counterpart of the trajectory trace for wave dynamical systems and can be computed based on the wave function and its derivative [55, 56, 59, 69, 7], see Eq. (10) below.
The 2D Husimi function is based on $\psi(s)$ and $\psi^{\prime}(s)$ where $s$ is the coordinate along the system boundary and indicates the normal derivative. The other phase space coordinate is $\sin\chi$, the sine of the (classical) angle of incidence that is related to angular momentum $m$ by $\sin\chi=m/n\mathrm{Re}\,(\Omega)$ with $\mathrm{Re}\,(\Omega)=kR_{0}$. The Husimi functions $H(s,\sin\chi)$ for the wave function in Figs. 1(a,b) are shown in Figs. 1(c,d). The mode is a radially symmetric whispering-gallery mode, and the angular momentum is thus conserved. Note that this is the simplest case, as for deformed cavities more intricate wave function (and thus phase space) emerges.
@@ III.1 Cylindrical cavities of finite height
$\Delta+\tilde{n}^{2}(\bm{r})k^{2}]\begin{Bmatrix}\bm{E}(\bm{r})\\ \bm{B}(\bm{r})\end{Bmatrix}=\bm{0}$ (4)
@@
$\begin{aligned} \bm{E}_{1,\parallel}&=\bm{E}_{2,\parallel}&n_{1}^{2}\bm{E}_{1,\perp}&=n_{2}^{2}\bm{E}_{2,\perp}\\ \bm{B}_{1,\parallel}&=\bm{B}_{2,\parallel}&\bm{B}_{1,\perp}&=\bm{B}_{2,\perp}\quad\end{aligned}$ (5)
@@
Applying the appropriate boundary conditions leads to a quantization in the $z$-direction, described by the quantum number $\eta\in\mathbb{N}_{0}$. So in general, modes are identified by their polarization and the three quantum numbers $(\rho,m,\eta)$. Examples for confinement in the $z$ direction for TM and TE modes are shown in Fig. 1(e) and (f). The TM mode is more confined in $z$ direction due to the discontinuity in $E_{z}$ described in Eq. (5).
@@ III.2 Extension of the Husimi formalism to three-dimensional cavities
For realistic 3D systems that we now describe, the new generalized coordinates $(\bm{q},\bm{p})$ with $\bm{q},\bm{p}\in\mathbb{R}^{3}$ result in a 6D phase space. It reduces to four dimensions after considering conservation of energy and projection to a surface of section that we choose to be the cylinder mantle here. We map the wave function onto the cylinder mantle using $\psi(\varphi,z)\rightarrow\psi(s,z)$, where $s=\varphi R_{0}$ is the arclength along the cavity surface. The resulting boundary conditions are periodic in $s$ with $\psi(s,z)=\psi(s\pm 2\pi R_{0},z)$. Here, we focus on the mantle as the main surface contribution, although top and bottom surfaces could be included straightforwardly when needed, e.g., for flat cavities.
Figure 2: Coherent state $\xi(\bm{q}^{\prime};\bm{q},\bm{p})$ on the mantle of a slanted, cylindrical cavity. It is measured at $\bm{q}^{\prime}$, where the value is determined by the geodesics $\bm{g}_{\ell},\bm{g}_{\mathrm{p}},\bm{g}_{\mathrm{o}}$ and the point $\bm{\tilde{q}}$.
@@
$\xi(\bm{q}^{\prime};\bm{q},\bm{p})=\frac{1}{\sqrt{\pi\sigma^{2}}}\times\\ \exp\left[-\mathcal{L}^{2}(\bm{q}^{\prime};\bm{q})/2\sigma^{2}-i\,\mathcal{P}(\bm{q}^{\prime};\bm{q},\bm{p})\right]$ (8)
with position $\bm{q}$, momentum $|\bm{p}|\leq nk$ on the cavity surface, and a Gaussian spatial profile characterized by $\sigma^{2}$. $\mathcal{L}$ is the geodetic distance between $\bm{q}$ and $\bm{q^{\prime}}$ along $\bm{g}_{\ell}$ (see Fig. 2). The second term is related to the coherent state’s angular momentum that is encoded in the phase $\mathcal{P}$ of the coherent state. The phase at $\bm{q}^{\prime}$ is obtained by projection on $\bm{g}_{\mathrm{p}}$ via a geodesic $\bm{g}_{\mathrm{o}}$ orthogonal to $\bm{g}_{\mathrm{p}}$. We call the projected point $\bm{\tilde{q}}$, where the phase is given as $\mathcal{P}(\bm{\tilde{q}};\bm{q},\bm{p})=|\bm{p}|\,\mathcal{L}(\bm{\tilde{q}};\bm{q})=\mathcal{P}(\bm{q}^{\prime};\bm{q},\bm{p})$ that reduces to the well-known scalar product form of the phase term, $\bm{p}\,\bm{r_{\bm{q},\bm{q^{\prime}}}}$ (with $\bm{r_{\bm{q},\bm{q^{\prime}}}}$ the vector from $\bm{q}$ to $\bm{q^{\prime}}$) in euclidean geometries.
We then define the overlap of the coherent state with the wave function $\psi$ at the boundary as
@@
with angular momentum weight $\mathcal{F}(p)=\sqrt{n}(1-\sin^{2}\chi_{0})^{1/4}$.
For cylindrical (or conic) cavities the geodesics can be obtained trivially, as all straight lines on the unrolled surface are geodesics. Then, $\mathcal{L}$ and $\mathcal{P}$ can be evaluated as if the cavity mantle were a flat surface using scalar products. For the phase space coordinates we follow the conventional notation of the Husimi function formalism [55]. For cylindrical 3D cavities the Husimi function $H=H(\bm{q},\bm{p})$ depends on the spatial coordinates $\bm{q}=(s,z)$ and on the momentum coordinates $\bm{p}=(nk\sin\chi_{xy},nk\sin\chi_{z})$, see Fig. 3(a), distinguishing between components orthogonal and parallel to the $z$-axis. The momentum coordinates can also be written in polar representation $(nk\sin\chi_{0},\phi_{xy})$ with total momentum in the surface (mantle) plane $n^{2}k^{2}\sin^{2}\chi_{0}=|\bm{p}|^{2}=n^{2}k^{2}(\sin^{2}\chi_{xy}+\sin^{2}\chi_{z})$ and angle $\phi_{xy}$ to the $xy$-plane where $\tan\phi_{xy}=\sin\chi_{z}/\sin\chi_{xy}$.
To illustrate this generalized Husimi function in the new phase space spanned by $\sin\chi_{xy}$ and $\sin\chi_{z}$, it is helpful to consider a simple test system: A wave function $\psi$ is given by a linear combination of plane waves, which propagate along a two-dimensional surface as shown in Fig. 3(b-c). The Husimi function is calculated for a representative, fixed position $\bm{q}$ while the momentum $\bm{p}$ is varied. A single plane wave can be identified trivially in real space and gives a single peak in phase space in Fig. 3(d). For two (or more) overlapping plane waves in Fig. 3(c) the field function can become arbitrarily complex, making it impossible to distinguish components in real space. The phase space recovers the constituent plane waves and their parameters as shown in Fig. 3(e).
@@ III.3 Husimi functions for cylindrical cavities
Figure 5: Field intensity (left) and Husimi functions (right) for TM modes, with $(\rho,m)=(0,6)$ and various $\eta$. The $\eta$ and frequencies are (a) $(\eta,\Omega)=(0,3.4402-i\,0.0049)$, (b) $(\eta,\Omega)=(3,3.5578-i\,0.0014)$ and (c) $(\eta,\Omega)=(5,3.7182-i\,0.0006)$. The phase space functions are shown for $\sin\chi_{xy}>0$, the rest is given by symmetry.
@@
The Husimi function for the $(\rho,m,\eta)=(0,5,0)$-TM and TE modes with $h/R=0.75$ are shown in Fig. 1(e) and (f). The phase space figures are calculated by using a constant coherent state position $\bm{q}=\mathrm{const}$ and varying the momentum coordinates $\bm{p}$. This gives a circular border for the parameter space from the constraint $\sin^{2}\chi_{xy}+\sin^{2}\chi_{z}<1$. Due to the rotational symmetry, only coherent state positions with angles $\varphi\in[0,\pi/2m]$ need to be considered. Given the mirror symmetry at $z=0$, we focus on coherent states with $z\in[0,-h/2]$ without loss of generality.
@@
A sharp peak is present at $\sin\chi_{xy}=\pm m/n\,\mathrm{Re}\,\Omega$ for all values of $\sin\chi_{z}$. This is analogous to the result of the 2D cavity in Figs. 1(c,f). The phase space is mirrored along $\sin\chi_{xy}=0$ because the standing wave considered here is a superposition of two traveling waves with $\pm m$. In addition, the existence of the geometry symmetry plane, $z=0$, results in a phase space mirror symmetry with axis $\sin\chi_{z}=0$. For the TE mode, the intensity maximum is closer to the origin which reflects its shorter lifetime since a larger part of the phase space intensity is located in the leaky region.
@@
Another degree of freedom we need to take into account is the localization of the coherent state, given by its standard deviation $\sigma$ in Eq. (8). Fig. 4 compares the phase spaces for two values of $\sigma$ for various points $\bm{q}_{i}$ on the cylinder surface. Figures 4(b,d,f) and (h) show the results for strong localization with $\sigma/R_{0}=0.32$, corresponding to a full-width at half-maximum of $\lambda/n$, where $\lambda$ is the vacuum wavelength. The intensity distribution is different between $\bm{q}_{1}$ and $\bm{q}_{2}$ in Figs. 4(b) and (d). $\bm{q}_{1}$ is located on an intensity maximum, leading to $H(\bm{q}=\bm{q}_{1},\bm{p}=\bm{0})>0$. The nodal line at $\bm{q}_{2}$ results in $H(\bm{q}=\bm{q}_{2},\bm{p}=\bm{0})=0$, since the fields of two neighboring intensity maxima average to zero due to their different signs. The Husimi functions in Figs. 4(f,h) are taken at points closer to the cavity boundary. The intensity at $\bm{q}=\bm{q}_{3},\bm{q}_{4}$ is significantly lower than at $\bm{q}=\bm{q}_{1},\bm{q}_{2}$, which is due to the lower overall field intensity at the cylinder end faces. Spatial dependence is very pronounced for localized coherent states and is lost for the weakly localized state with $\sigma=0.95$, see Figs. 4(c,e,g) and i.
@@
Fig. 5 explores the dependence of the Husimi function on the quantization in $z$ for a cylinder with $h/R=5$. It shows the behavior of the Husimi function with $\sigma/R_{0}=0.95$ for $\eta=0,3,5$ for TM modes with $(\rho,m)=(0,6)$. The wave functions are now described by sharply localized intensity maxima, in contrast to the previously spread out maxima in Fig. 4. When $\eta$ is increased in Figs. 5(b) and (c), the angle $\varphi_{xy}$ of the maximum increases. This is related to the distribution of wavenumber components in the in- and out-of-plane components. The mode performs whispering-gallery type motion in $xy$ and a ”bouncing ball” type motion in $z$. The rays can be thought of as angled at $\tan\phi_{xy}=k_{z}/k_{xy}$ (compare Fig. 3(a)). Increasing $\eta$ results in more energy being stored in the bouncing ball motion, thereby increasing $\sin\chi_{z}$ of the intensity maximum. The symmetry in the $xy$-plane results in two symmetry-related signatures at $\pm\sin\chi_{z}$. The wave function on the cylinder mantle displays a ”checkerboard” pattern, which can be recovered by the superposition of two plane waves at angles $\pm\varphi_{xy}$ (compare Figs. 3(c,e)). For TE modes the behavior is similar, which is demonstrated in Appendix A.